Find the length of the diagonal of a rectangle with l = 9 cm and w = 3 cm.
1
Expert's answer
2013-04-30T08:55:52-0400
According to the Pythagorean theorem, the length of the diagonal of a rectangle is

L = √(l² + w²) = √(9² + 3²) = √(81 + 9) = √90 = 3√10 ≈ 9.49 [cm].
Answer: 9.49 cm.